Campbell Supper Co. paid a $0.702 dividend per share in 2013, which grew to $0.86 in 2016. This growth is expected to continue. What is the value of this stock at the beginning of 2017 when the required return is 8.7 percent? (Round the growth rate, g, to 4 decimal places. Round your final answer to 2 decimal places.)
Answer #1

We use the formula:
A=P(1+r/100)^n
where
A=future value
P=present value
r=rate of interest
n=time period.

0.86=0.702*(1+r/100)^3

(0.86/0.702)^(1/3)=(1+r/100)

(1+r/100)=1.07

r=1.07-1

=0.07(Approx)

Hence Current price=D1/(Required return-Growth rate)

=(0.86*1.07)/(0.087-0.07)

=$54.13(Approx).